# Hardware test and validation of the angular droop control: Analysis and experiments

### v1

@@ Discussion
Figures 22 and 23 show the experimental results of the power sharing experiment. Assumption 1 is here again satisfied and that is seen from Figure 22, where $\lVert\theta^{s}_{1}-\theta^{s}_{2}\rVert<0.1$. For $\gamma_{1}=\gamma_{2}=500$, power sharing is guaranteed at steady state, where $P^{s}_{k}\approx P^{*}/2$ for $k=1,2$. This corresponds to zero reactive power at steady state. As expected, the two converters synchronize in frequency and their angles converge respectively to frequency synchronous steady states within 0.50.50.5\,0.5s. The trade-off between active power sharing and the induced steady state modulation angle is visible when comparing Figure 18 and 22. Thus, our experiment validates the power sharing capabilities of the angular-droop controlled DC/AC converter system, one of the most important plug and play properties for converter control design in power networks.
VII Conclusion
We validated the grid-forming capabilities of the angular droop control through two test scenarios embedded within the hardware experimentation in the Smart Energy System Control Laboratory at KIT. In both scenarios, the DC/AC converter in closed-loop with the angular droop control recovers nominal operation following a blackout, provides active power, also upon a sudden change in the load, achieves frequency synchronization and power sharing in a network setting upon suitable tuning. Our future work aims at addressing the compatibility of the angular with frequency-droop control and therefore the interoperability with the remainder of the grid components. Another venue is concerned with voltage control and large-scale hardware experimentation.

### v2

@@ Discussion
Figure 20 shows the experimental results of the power-sharing experiment. The security constraint in [17] is here again satisfied with $\lVert\theta^{s}_{1}-\theta^{s}_{2}\rVert<0.1$. For $\gamma_{1}=\gamma_{2}=500$, power-sharing is guaranteed at steady state, where $P^{s}_{k}\approx P^{*}/2$ for $k=1,2$. This corresponds to zero reactive power at steady state. As expected, the two converters synchronize in frequency and their angles converge respectively to frequency synchronous steady states within 0.50.5\,s. Thus, our experiment validates the power-sharing capabilities of the angular-droop controlled DC/AC converter system, one of the most important plug and play properties for converter control design in power networks.
V Conclusion
We demonstrate the grid-forming properties of the angular droop control in two scenarios embedded in a hardware experiment setup. For this we provide traceable analysis and solutions to the challenges that arise from virtual angle control in real-world applications. First, the discretized control law is rewritten to conform with the restrictions posed by hardware implementation. Second, the susceptibility to clock drifts in multi-converter settings is analyzed and resolved by distributing a common master clock. Using a single converter setup, the black start capabilities of angular droop control as well as the capability to withstand sudden load changes while returning to zero frequency deviation is shown. With the extension to a multi-converter setup the frequency synchronization as well as load sharing properties are demonstrated and tuning guidelines are established. Our future work aims to extend the angular droop control with voltage regulation, therefore relaxing the assumption on constant AC voltage amplitude. Furthermore, the interoperability of angular droop with other control strategies needs to be considered. For prospective industrial-microgrid applications, the influence of nonlinear loads as well as the extension to larger multi-converter setup will be investigated.
Author Contributions
Taouba Jouini: conceptualization, data curation, formal analysis, investigation, methodology, visualization, writing - original draft, writing - review & editing Jan Wachter: conceptualization, data curation, formal analysis, investigation, methodology, visualization, writing - original draft, writing - review & editing Sophie An: conceptualization, formal analysis, investigation, methodology, writing - original draft Veit Hagenmeyer: conceptualization, funding acquisition, supervision, writing - review & editing
Conflict of Interest
The authors declare no conflict of interest.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A High-order converter modeling and control
This section details how the angular droop control law (2) is implemented on the experiment hardware. With some abuse of notation, we omit throughout subsequent sections the subscript $k$ to denote a quantity $z_{k}$ of the $k$-th converter.
A-A Modeling and control DC/AC converter
We start by relaxing our modeling assumptions from Section II towards a hardware experiment in two directions:
  • First, we include a DC power supply and a DC/DC converter relying on vector control consisting of cascaded voltage and current control loops put in series behind the DC/AC converter to provide adequate DC-link voltage.
Hereby, we further assume that all AC voltage amplitudes are constant and at nominal.
A-A1 Modeling DC/AC converter
We consider a three-phase, averaged and balanced DC/AC converter as shown in Fig. 2 and described in abcabc-frame [24, 21],
$L\,\dot{i}$ $=-R\,i+\frac{1}{2}\overline{u}\,\,V_{dc}-v,$ (36)
$C\,\dot{v}$ $=-G\,v+i-i_{o},$
where $V_{dc}$ denotes the DC side capacitor voltage. On the AC side, let $i\in^{3}$ denote the inductance current and $v\in^{3}$ the output voltage. The filter resistance and inductance are specified by $R>0$ and $L>0$, respectively. The capacitor $C>0$ is set in parallel with the load conductance $G>0$ to ground. The DC/AC converter is connected to the AC network, where $i_{o}\in$ is the output current flowing into the network. Note that the pulse width modulation signal $\overline{u}\in[-1,1]$ relates to the converter duty cycle $d_{AC}\in[0,1]$ via
$d_{AC}=\frac{1}{2}+\frac{\overline{u}}{2},$ (37)
where $\overline{u}$ represents the main input to the DC/AC converter.
A-A2 Direct implementation
First, we define the active power $P=v^{\top}i_{o}$, and the nominal steady state $P^{*}=v^{*\top}i_{o}^{*}$. For the direct implementation of the angular droop control we propose
where $0<A<1$ is the amplitude of the modulation signal $\overline{u}$. Fig. 21 depicts a summarizing block diagram of the direct implementation of the angular droop control (2). Note that in (38), the angular droop control increments the converter’s state with a virtual angle $\theta$ through the modulation signal $\bar{u}$.
[…]
Figure 21: Implementation of the angular droop for the DC/AC converter via direct control of the modulation signal $\bar{u}$.
A-A3 Indirect Implementation
We propose an indirect implementation of the angular droop control that relies on cascaded control. In particular, the indirect implementation entails inner voltage and current control loop according to a cascaded architecture. After a Park transformation $\mathcal{P}(\theta_{dq})$ with angle $\theta_{dq}(t):=\theta(t)$ and given the reference voltage $v^{d}$ in abcabc-frame [24],
$i_{dq}^{d}=Y\,v_{dq}+i_{o,dq}-k_{VP}(v_{dq}-v_{dq}^{d})-k_{VI}\int_{0}^{t}\big(v_{dq}(\tau)-v_{dq}^{d}\big)\mathrm{d}\tau,$ (40)
where $k_{VP},k_{VI}~>0$ are the control gains and $Y=G+C\mathbf{J}\omega^{*}$. To track the reference current (40), we design an inner current loop based on PI control using the switching voltage $v^{d}_{m}=\frac{1}{2}\overline{u}_{dq}\,V_{dc}$ as follows
$v_{m}^{d}=Z\,i_{dq}+v_{dq}-k_{IP}(i_{dq}-i^{d}_{dq})-k_{II}\int_{0}^{t}(i_{dq}(\tau)-i_{dq}^{d})\,\mathrm{d}\tau,$ (41)
where $k_{IP},k_{II}>0$ are control gains and $Z=R+L\,\mathbf{J}\,\omega^{*}$. By applying the inverse Park transformation $\mathcal{P}^{-1}(\theta_{dq})$, we recover the modulation input $\overline{u}$ in abcabc-frame as follows,
It is noteworthy that, the pairs $(k_{VP},k_{VI})$ and $(k_{IP},k_{II})$ are chosen to guarantee time-scale separation, where the current control loop is faster than the voltage control loop. Fig. 22 summarizes the overall scheme of the indirect implementation based on cascaded control.
[…]
Figure 22: Indirect cascaded control of the angular droop after dqdq-transformation of the DC/AC converter. PI-VC and PI-$C$ denote the proportional integral (PI) voltage and current controller, respectively.
A-B Modeling and control DC/DC boost converter
A-B1 Modeling DC/DC boost converter
Figure 23: A schematic representation of the boost converter consisting of a half-bridge module and an inductance $L_{b}$ with the parasitic resistance $R_{b}$.
$L_{b}\dot{I}_{b}$ $=-R_{b}\,I_{b}+V_{b}-V_{c}\,$, (43)
$C_{dc}\dot{V}_{dc}$ $=-G_{dc}V_{dc}+I_{dc},\quad I_{dc}=\frac{V_{b}I_{b}}{V_{dc}}\,.$
$V_{c}=(1-d)\cdot V_{dc}\in.$ (44)
A-B2 Boost converter
[…]
The cascaded control architecture to regulate the DC-bus voltage of the boost converter exploits the differential equations (43) and is summarized in Fig. 24. In particular, an outer loop regulates the DC capacitor voltage $V_{dc}$ at a nominal value $V_{dc}^{*}>V_{b}>0$ by specifying a reference current $I^{d}_{b}\in$ given explicitly by,
$d=1-\frac{V_{b}-V^{d}_{l}}{V_{dc}}.$ (47)
[Uncaptioned image] Taouba Jouini received her PhD degree in System and Control Engineering from Lund University in 2022. She is currently a postdoctoral researcher at Leibniz University of Hannover. Her research interests are centered around the optimal management of multi-modal energy systems and the development of test environments for autonomous vehicles.
[Uncaptioned image] Jan Wachter received the M.Sc. degree in mechanical engineering from the Karlsruhe Institute of Technology (KIT), Karlsruhe, Germany. He is currently working towards the Ph.d. degree at the Institute for Automation and Applied Informatics (IAI) of the KIT. His research interest includes the control of grid connected converters systems and the challenges arising from their interaction.
[Uncaptioned image] Sophie Xing An received the M.Sc. degree in Electrical Engineering and Information Technology from the Karlsruhe Institute of Technology (KIT), Karlsruhe, Germany. She was with the Institute for Automation and Applied Informatics (IAI) of the KIT. Her research interest includes the control of the power grid in the face of the energy transition.
[Uncaptioned image] Veit Hagenmeyer received the Ph.D. degree from Université Paris XI, Paris, France, in 2002. He is currently a Professor of energy informatics with the Faculty of Computer Science, and also the Director with the Institute for Automation and Applied Informatics (IAI), Karlsruhe Institute of Technology, Karlsruhe, Germany. His research interests include modeling, optimization and control of sector-integrated energy systems.